NvidiaIntely (nvidiaintely):

Find the height of a square pyramid with volume 104.06 ft3 and dimensions of base are 5 ft by 5 ft. 25 ft 12.5 ft 8.325 ft 6.25 ft

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (anglina4196):

Is it ok if I help?

NvidiaIntely (nvidiaintely):

Yes, but please be fast. @Kevin, you try also :)

OpenStudy (anglina4196):

ok

OpenStudy (kevin):

104.06 = 1/3*5*5*t t = 12.5

OpenStudy (kevin):

@anglina pls switch me after this. I'm gonna off :D

OpenStudy (anglina4196):

The equation for the volume of a pyramid is V = 1/3b*h If you're given the volume and the area of the base, you can solve for the height 104.06 = 1/3 (25) * h I just plugged in the volume and 25 for the base since it's 5 x 5. From here all you have to do is solve for h. Hope this helps.

OpenStudy (anglina4196):

ok